### v1

# LLM-Based Multi-Agent Systems for Software Engineering: Vision and the Road Ahead

Abstract.
Integrating Large Language Models (LLMs) into autonomous agents marks a significant shift in the research landscape by offering cognitive abilities competitive to human planning and reasoning. This paper envisions the evolution of LLM-based Multi-Agent (LMA) systems in addressing complex and multi-faceted software engineering challenges. LMA systems introduce numerous benefits, including enhanced robustness through collaborative cross-examination, autonomous problem-solving, and scalable solutions to complex software projects. By examining the role of LMA systems in future software engineering practices, this vision paper highlights the potential applications and emerging challenges. We further point to specific opportunities for research and conclude with a research agenda with a set of research questions to guide future research directions.
Large Language Models, Autonomous Agents, Multi-Agent Systems, Software Engineering
conference: International Workshop on Software Engineering in 2030; November 2024; Porto de Galinhas (Brazil)
1. Introduction
Autonomous agents, defined as intelligent entities that autonomously perform specific tasks through environmental perception, strategic self-planning, and action execution (Franklin and Graesser, 1996; Albrecht and Stone, 2018; Mele, 2001), have emerged as a rapidly expanding research field since the 1990s (Maes, 1993). Despite initial advancements, these early iterations often lack the sophistication of human intelligence (Unland, 2015). However, the recent advent of Large Language Models (LLMs) (Kasneci et al., 2023) has marked a turning point. This LLM breakthrough has demonstrated cognitive abilities nearing human levels in planning and reasoning (Achiam et al., 2023; Kasneci et al., 2023), which aligns with the expectations for autonomous agents. As a result, there is an increased research interest in integrating LLMs at the core of autonomous agents (Lo, 2023; Xi et al., 2023; Wang et al., 2023b) (for short, we refer to them as LLM-based agents in this paper).
Nevertheless, the application of singular LLM-based agents encounters limitations, since real-world problems often span multiple domains, requiring expertise from various fields. In response to this challenge, developing LLM-Based Multi-Agent (LMA) systems represents a pivotal evolution, aiming to boost performance via synergistic collaboration. An LMA system harnesses the strengths of multiple specialized agents, each with unique skills and responsibilities. These agents work in concert towards a common goal, engaging in collaborative activities like debate and discussion. These collaborative mechanisms have been proven to be instrumental in encouraging divergent thinking (Liang et al., 2023), enhancing factuality and reasoning (Du et al., 2023), and ensuring thorough validation (Wu et al., 2023b). As a result, LMA systems hold promise in addressing a wide range of complicated real-world scenarios across various sectors (Horton, 2023; Wang et al., 2023c, a), such as software engineering (Lo, 2023; Qian et al., 2023; Li et al., 2024; Hong et al., 2023).
The study of software engineering (SE) focuses on the entire lifecycle of software systems (Kan, 2003), including stages like requirements elicitation (Goguen and Linde, 1993), development (Abrahamsson et al., 2017), and quality assurance (Tian, 2005), among others. This multifaceted discipline requires a broad spectrum of knowledge and skills to effectively tackle its inherent challenges in each stage. Integrating LMA systems into software engineering introduces numerous benefits:
  1. (1)
    Robustness and Fault Tolerance: LMA systems address robustness issues through cross-examination in decision-making, akin to code reviews and automated testing frameworks, thus detecting and correcting faults early in the development process. On their own, LLMs may produce unreliable outputs, known as hallucination (Zhang et al., 2023; Yang et al., 2023), which can lead to bugs or system failure in software development. However, by employing methods like debating, examining, or validating responses from multiple agents, LMA systems ensure convergence on a single, more accurate and robust solution. This enhances the system’s reliability, error handling, and aligns with best practices in software quality assurance.
  2. (2)
    Autonomous Problem-Solving: LMA systems can bring significant autonomy to SE tasks. It is an intuitive approach to divide high-level requirements into sub-tasks and detailed implementation, which mirrors agile and iterative methodologies (Larman, 2004) where tasks are broken down and assigned to specialized teams or individuals. By automating this process, developers are freed to focus on strategic planning, design thinking, and innovation.
  3. (3)
    Scalability to Complex Systems: The growth in complexity of software systems, with increasing lines of code, frameworks, and interdependencies, demands scalable solutions in project management and development practices. LMA systems offer an effective scaling solution by incorporating additional agents for new technologies and reallocating tasks among agents based on evolving project needs. LMA systems ensure that complex projects, which may be overwhelming for individual developers or traditional teams, can be managed effectively through distributed intelligence and collaborative agent frameworks.
Existing research has illuminated the critical roles of these collaborative agents in advancing toward the era of Software Engineering 2.0 (Lo, 2023). LMA systems are expected to significantly speed up software development and drive innovation, marking a major shift in software engineering practices. This article aims to delve deeper into the roles of LMA systems in shaping the future of software engineering, spotlighting potential applications, emerging challenges, and the vital collaboration between academia and industry to harness these advancements fully. Furthermore, we present a comprehensive research agenda, which enumerates a series of research questions to direct subsequent research efforts.
2. Preliminary
2.1. Autonomous Agent
An autonomous agent is a computational entity designed for independent and effective operation in dynamic environments (Mele, 2001). Its essential attributes are:
  • Autonomy: Independently manages its actions and internal state without external controls.
  • Perception: Detects the changes in the surrounding environment through sensory mechanisms.
  • Intelligence and Goal-Driven: Aims for specific goals using domain-specific knowledge and problem-solving abilities.
  • Social Ability: Can interact with humans or other agents, manages social relationships to achieve goals.
  • Learning Capabilities: Continuously adapts, learns, and integrates new knowledge and experiences.
2.2. LLM-based Autonomous Agent
Formally speaking, an LLM-based agent can be described by the tuple $\langle L,O,M,A,R\rangle$ (Cheng et al., 2024), where:
  • $L$ symbolizes the Large Language Model, serving as the agent’s cognitive core. It is usually equipped with extensive, potentially domain-specific, knowledge, enabling it to make informed decisions from observations, feedback, and rewards. Typically, an LLM possesses significant zero-shot or few-shot learning capabilities, and we interact with the LLM with prompts.
  • $O$ stands for the Objective, the desired outcome or goal the agent aims to achieve. This defines the agent’s focus, driving its strategic planning and task breakdown.
  • $M$ represents Memory, which holds information on both historical and current states, as well as feedback from external interactions.
  • $A$ signifies Action, encompassing the range of executions of the agent, from utilizing tools to communicating with other agents.
  • $R$ refers to Rethink, a post-action reflective thinking process that evaluates the results and feedback, along with stored memories. Guided by this insight, the LLM-based agent then takes subsequent actions.
2.3. LLM-Based Multi-Agent Systems
In contrast to single-agent systems, a Multi-Agent system places a greater emphasis on coordination, communication, and knowledge sharing among agents. LMA systems are especially suited for complex tasks that require expertise across various knowledge domains. In this paper, we define that a Multi-Agent system, particularly an LLM-Based Multi-Agent system, includes two primary elements: an orchestration platform and LLM-based agents.
Orchestration Platform: The platform serves as the core infrastructure that manages interaction and information flow among agents. It is designed to ensure efficient operation and communication within the LMA system through several key functions: Agent Profiling: As we desire a broad spectrum of expertise, agents are defined with unique abilities and actions, fitting specific roles within the system. Profiles may be explicitly pre-defined (Qian et al., 2023; Hong et al., 2023) or dynamically generated by LLMs (Wang et al., 2023b).
Communication Mechanisms: Specifically, these communication mechanisms include:
  • Communication Paradigms: Styles of interaction, such as Cooperative (i.e., agents work together towards a shared goal) (Abdelnabi et al., 2023), Debate (i.e., agents engage in discussions to challenge and refine ideas) (Li et al., 2024), and Competitive (i.e., agents have conflicting objectives and compete against each other) (Wu et al., 2024).
  • Communication Structure: The organization of communication channels, which can be:
      Decentralized: Agents communicate directly with each other without a central controlling authority (Chen et al., 2023).
      Centralized: A central agent facilitates communication between all other agents (Agashe, 2023).
      Hierarchical: With clear layers of authority and responsibility, information flows from higher to lower levels (Zhao et al., 2024).
      Nested: Communication structures are embedded within one another, combining characteristics of hierarchical, decentralized, and centralized systems (Han et al., 2024).
  • Communication Content: The actual data exchanged among agents, often in text form. In the context of software engineering, this may be programming languages, commit messages (Zhou et al., 2023), Stack Overflow posts (He et al., 2022), etc.
LLM-based agents: We focus on the scenarios where agents within the Multi-Agent systems are powered with LLMs. Leveraging LLMs’ feedback integration capabilities, agents can engage in meaningful dialogues, enhancing cooperation and performance.
3. Current Development
In this section, we review the current applications of LMA agents in the software engineering domains.
Software Development. Several works have applied LMA systems in software development. Li et al. introduce a pioneering communicative Multi-Agent framework named role-playing (Li et al., 2024) with exceptional performance in code generation. ChatDev (Qian et al., 2023) divides the development process into designing, coding, testing, and documentation phases, with specialized teams of software agents, such as programmers and test engineers. Experiment results demonstrate that ChatDev can finalize the development of an application in less than seven minutes with a cost below one dollar. MetaGPT (Hong et al., 2023) embeds Standardized Operating Procedures (SOPs) (Belbin and Brown, 2022) into its workflow and further includes roles like product manager and QA Engineer. Subsequent studies have further explored the impact of role differentiation among LLM-based agents on collaborative software development  (Dong et al., 2023; Huang et al., 2023).
Frameworks. Numerous open-source repositories have been developed, which offer developers the means to customize and refine these agents for distinct tasks. AutoGen (Wu et al., 2023a) is an open-source framework that enables the creation of LMA applications through customizable, conversable agents that integrate LLMs, human inputs, and executable tools with flexible conversation patterns. Langroid (langroid contributors, 2024) implements its Multi-Agent pipeline based on the Actor Framework (Hewitt et al., 1973). Moreover, frameworks like LangChain (Chase, 2022), OpenAgents (Xie et al., 2023), AutoGPT (Significant Gravitas, [n. d.]) and GPT-Engineer (Osika, 2023), can also be adaptable for use in LMA systems, despite being originally designed on single-agent applications.
4. Research Agenda
Previous research has laid the groundwork for the exploration of LMA systems in software engineering, yet this domain remains in its nascent stages, with many critical challenges awaiting resolution. In this section, we outline our perspective on these challenges and suggest research questions that could advance this burgeoning field. We envision two phases for the development of LMA in software engineering. We discuss each of these phases below and suggest a series of research questions that could form the basis of future research projects.
Phase 1: Enhancing Individual Agent Capabilities
Indeed, the effectiveness of an LMA system is closely linked to the capabilities of its individual agents. This first phase is dedicated to improving these agents’ skills, with a particular focus on adaptability and the acquisition of specialized skills in SE. The potential of individual LLM-based agents in SE is further explored through our initial research questions:
  1. (1)
    What SE roles are suitable for LLM-based agents to play and how can their abilities be enhanced to represent these roles?
  2. (2)
    How to design an effective, flexible, and robust prompting language that enhances LLM-based agents’ capabilities?
Refining Role-playing Capabilities in SE. The capability of the LLM-based agents for role-playing is pivotal under an LMA system. These agents demonstrate proficiency in simulating well-recognized roles, but they may not accurately represent specialized SE roles, such as blockchain developers and DevOps Engineers. This shortfall arises from their training on general text corpora, which typically lack the specialized content necessary to grasp the nuanced expertise required in SE. To address this limitation, we propose a structured three-step approach encompassing the identification, selection, and enhancement of role-playing abilities.
The initial step focuses on identifying key SE roles, prioritizing those with high industry demand, significant complexity, and the potential to substantially boost productivity. This involves a detailed analysis of the SE sector’s current trends and needs, utilizing market reports, job postings, and industry forecasts to identify roles where LLM-based agents can offer the greatest value. The process concludes with value addition modeling, which evaluates potential efficiency gains, cost reductions, and the acceleration of innovation, informing the selection of roles for LLM-based agents.
Progressing to the second step, the focus shifts towards understanding the limitations of LLM-based agents against the demands of identified SE roles. Through competency mapping, this phase evaluates the agents’ proficiency in essential skills such as code generation, architecture design, and project management. By employing real-world software engineering tasks, we can accurately assess the agents’ performance and identify specific areas where their skills may be lacking. Furthermore, consultations with software engineering professionals and academics are instrumental in refining the capabilities of LLM-based agents. These discussions uncover the complex needs of specialized roles that training data may not fully represent, guiding targeted improvements in the agents’ skill sets.
The last step is to tailor the LLM-based agents’ to effectively represent the identified SE roles. This requires creating specialized training datasets that reflect the unique requirements of each role for fine-tuning. Specifically, it involves collecting materials from open-source code repositories, technical forums, and detailed software documentation to build a comprehensive corpus for training agents. Additionally, we can design customized prompts that enhance the adaptability of LLM-based agents to specific SE roles.
Advancing Prompts through Agent-oriented Programming Paradigms
In previous sections, we highlight that well-crafted prompts can substantially enhance an LLM-based agent’s performance, including the agents’ abilities in planning, reasoning, communication, and mitigating the issue of hallucinations. Yet, crafting prompts for LLM-based agents is particularly challenging. The complexity lies in developing a versatile, effective, and robust prompt framework that is suitable across various scenarios and roles. Currently, this area of research is still largely unexplored and urgently requires the development of an advanced prompting language designed to augment the cognitive functions of LLM-based agents. Introducing Agent-Oriented Programming (AOP) (Shoham, 1993) into the realm of LLM-based agents holds promising potential. Similar to objects in Object-Oriented Programming (OOP) (Wegner, 1990), AOP is a paradigm that treats agents as fundamental programming units. This approach emphasizes defining agents’ reasoning mechanisms, objectives, and methods of cooperation. By adopting AOP, we can craft a more refined prompting language, one that facilitates the articulation of complex, detailed instructions alongside specific constraints, This, in turn, will empower LLM-based agents to execute their designated roles with heightened efficiency and precision.
Phase Two: Optimizing Agent Synergy
In Phase Two, the spotlight turns towards optimizing agent synergy, underscoring the importance of collaboration and how to leverage the diverse strengths of individual agents. This phase delves into both the internal dynamics among agents and the role of external human intervention in enhancing the efficacy of the LMA system. Key research questions guiding this phase include:
  1. (3)
    How to best allocate tasks between humans and LLM-based agents?
  2. (4)
    How to scale LMA systems for large-scale projects?
  3. (5)
    What industrial organization mechanisms can be applied to LMA systems?
  4. (6)
    What strategies allow LMA systems to dynamically adjust their approach?
  5. (7)
    How to ensure security among private data sharing within LMA systems?
Human-Agent Collaboration Strategically allocating tasks between humans and LMA agents to exploit their distinct strengths is crucial. Humans bring to the table unparalleled creativity, critical thinking, and ethical insight. Conversely, LMA agents are adept at swiftly processing extensive data sets, executing repetitive tasks with exceptional precision, and uncovering patterns that might escape human notice. Identifying the most effective roles for humans, optimizing the provision of human feedback, and determining the appropriate moments for human intervention in agent collaborations pose significant challenges. Furthermore, developing predictive models to ascertain the ideal human-to-agent ratio across various project types and stages is another fundamental concern. Such models need to evaluate project complexity, time restrictions, project priorities, and the particular capabilities and limitations of both human participants and LMA agents. As a result, it ensures that tasks are allocated to fully harness both human creativity and agent efficiency within the project.
Scaling Up for Complex Projects. Current applications of LMA systems are limited to small-scale software projects (Qian et al., 2023; Hong et al., 2023), however, transitioning to complex projects introduces several challenges: First, as software projects grow in size and complexity, breaking down high-level requirements into manageable sub-tasks for agents becomes increasingly difficult. Global planning and task allocation become more complicated since the variability and randomness in tasks escalate. Second, expanding projects necessitate a larger team of agents. This increase complicates the communication network, introducing challenges in coordination among agents. Such issues can lead to misinformation, threatening the resultant outcome of the LMA system. Furthermore, extensive information exchange strains memory and computational resources. A larger team of agents also implies more rounds of discussion are needed, which can significantly prolong the decision-making process. To address these challenges, innovative solutions in architectural design are required. These solutions should focus on facilitating efficient and reliable information exchange within the expanded LMA systems, ensuring the timely and effective completion of tasks. Streamlining communication and coordination among a growing number of agents, while managing the increased computational demands, are key to scaling LMA systems for complex software projects effectively.
Industrial Inspiration. The organizational strategies employed by industrial companies in structuring their development teams offer valuable lessons for LMA system design. Just as these companies adjust team sizes, roles, and project management approaches in response to shifting project landscapes, LMA systems can adopt similar adaptive mechanisms to remain agile and effective. For example, we can draw inspiration from agile development practices, and emulating iterative development models. LMA systems can implement mechanisms for incremental changes and continuous learning, allowing for gradual adaptation without overhauling the entire system.
Dynamic Adaptation. Given the complexity and variability of different software tasks, predicting the optimal configuration for LMA systems at the task’s outset is unrealistic. It therefore demands the system to dynamically adjust its scale and strategies while developing. This includes the ability to automatically modify the number of agents, redefine their roles, reallocate memory and computational resources, and modify communication methods in response to the development progress. Furthermore, as software projects inevitably change requirements, the adaptability of LMA systems to these changes becomes crucial. Through evaluating existing solutions and identifying reusable elements, LMA systems should be able to minimize redundant work. LMA systems should continuously learn from each cycle of development, identifying patterns, efficiencies, and inefficiencies. This learning enables the system to make informed decisions about how to approach similar tasks in the future or adapt existing solutions to meet new requirements. Moreover, the inherent uncertainty in the software development process poses another challenge in defining effective termination conditions of LMA systems. Relying solely on pre-defined termination conditions may not always be feasible, as these conditions might lead to infinite loops or premature task termination, hindering potential improvements. To address this, another critical research direction is the dynamic control of task termination. Possible solutions include implementing a monitoring model that continuously assesses the progress of tasks against their objectives. These models dynamically adjust their criteria for task completion based on real-time analysis, and may also escalate for human intervention judiciously, allowing for continuous refinement and optimization of outcomes.
Privacy and Partial Information. In multi-organization projects, data often resides in silos, leading to the need for specific agents to access sensitive information while ensuring privacy. This necessitates robust access control mechanisms that prevent unauthorized access but also address the varied data access needs among agents. Establishing protocols for agents to share insights from sensitive data without revealing the data itself is essential. Research could focus on applying differential privacy (Dwork, 2006) and secure multi-party computation (Goldreich, 1998) to enable involvement in collective decision-making without breaching privacy. Additionally, investigating integrated data storage for non-sensitive information could help minimize redundancy and achieve system-wide consistency.
5. Conclusion and Future Work
In conclusion, this paper outlines a structured research agenda to advance LMA integration in SE. Future work will concentrate on addressing the set of critical research questions identified, with a focus on enhancing LMA capabilities and optimizing their synergy within software development processes.
In the immediate term, efforts will be dedicated to enhancing the capabilities of LLM-based agents in representing specialized SE roles. This will involve creating specialized datasets and pre-training tasks that mirror the complex realities of SE tasks. Additionally, there will be a focus on formulating advanced prompting strategies, which can refine the agents’ cognitive functions and decision-making skills. Looking towards the longer-term objectives, the emphasis will transition towards optimizing the synergy between agents. The initial step involves examining optimal strategies for task allocation between humans and LLM-based agents, capitalizing on the unique strengths of both entities. Further, we need to delve into developing scalable methodologies for LMA systems, enabling LMA systems to efficiently orchestrate and complete large-scale, multifaceted SE projects. Furthermore, ensuring the privacy and confidentiality of data within LMA systems is critical. This entails exploring data management and access control mechanisms to protect sensitive information while still enabling the essential exchange of insights among project stakeholders. By systematically exploring these research questions, our goal is to drive innovation in LMA systems for software engineering, aiming to create a more cohesive, effective, and flexible LMA-driven development process.