### v2

# LLM-Based Multi-Agent Systems for Software Engineering: Vision and the Road Ahead

Abstract.
Integrating Large Language Models (LLMs) into autonomous agents marks a significant shift in the research landscape by offering cognitive abilities competitive to human planning and reasoning. This paper explores the transformative potential of integrating Large Language Models into Multi-Agent (LMA) systems for addressing complex challenges in software engineering. LLMs enhance agents’ cognitive abilities, enabling more robust collaboration, autonomous problem-solving, and scalable project management. We review the current application of LMA systems in software engineering and discuss how LMA systems can improve software development through cross-examination, error reduction, and task automation. Additionally, the paper outlines the emerging challenges and presents a research agenda aimed at advancing LMA system capabilities in software engineering, offering key research questions to guide future exploration.
Large Language Models, Autonomous Agents, Multi-Agent Systems, Software Engineering
1. Introduction
Autonomous agents, defined as intelligent entities that autonomously perform specific tasks through environmental perception, strategic self-planning, and action execution (Franklin and Graesser, 1996; Albrecht and Stone, 2018; Mele, 2001), have emerged as a rapidly expanding research field since the 1990s (Maes, 1993). Despite initial advancements, these early iterations often lack the sophistication of human intelligence (Unland, 2015). However, the recent advent of Large Language Models (LLMs) (Kasneci et al., 2023) has marked a turning point. This LLM breakthrough has demonstrated cognitive abilities nearing human levels in planning and reasoning (Achiam et al., 2023; Kasneci et al., 2023), which aligns with the expectations for autonomous agents. As a result, there is an increased research interest in integrating LLMs at the core of autonomous agents (Lo, 2023; Xi et al., 2023; Wang et al., 2023c) (for short, we refer to them as LLM-based agents in this paper).
Nevertheless, the application of singular LLM-based agents encounters limitations, since real-world problems often span multiple domains, requiring expertise from various fields. In response to this challenge, developing LLM-Based Multi-Agent (LMA) systems represents a pivotal evolution, aiming to boost performance via synergistic collaboration. An LMA system harnesses the strengths of multiple specialized agents, each with unique skills and responsibilities. These agents work in concert towards a common goal, engaging in collaborative activities like debate and discussion. These collaborative mechanisms have been proven to be instrumental in encouraging divergent thinking (Liang et al., 2023), enhancing factuality and reasoning (Du et al., 2023), and ensuring thorough validation (Wu et al., 2023). As a result, LMA systems hold promise in addressing a wide range of complicated real-world scenarios across various sectors (Horton, 2023; Wang et al., 2023e, a), such as software engineering (Lo, 2023; Qian et al., 2024c; Li et al., 2024; Hong et al., 2023).
The study of software engineering (SE) focuses on the entire lifecycle of software systems (Kan, 2003), including stages like requirements elicitation (Goguen and Linde, 1993), development (Abrahamsson et al., 2017), and quality assurance (Tian, 2005), among others. This multifaceted discipline requires a broad spectrum of knowledge and skills to effectively tackle its inherent challenges in each stage. Integrating LMA systems into software engineering introduces numerous benefits:
  1. (1)
    Robustness and Fault Tolerance: LMA systems address robustness issues through cross-examination in decision-making, akin to code reviews and automated testing frameworks, thus detecting and correcting faults early in the development process. On their own, LLMs may produce unreliable outputs, known as hallucination (Zhang et al., 2023; Yang et al., 2023b), which can lead to bugs or system failure in software development. However, by employing methods like debating, examining, or validating responses from multiple agents, LMA systems ensure convergence on a single, more accurate, and robust solution. This enhances the system’s reliability and aligns with best practices in software quality assurance.
  2. (2)
    Autonomous Problem-Solving: LMA systems can bring significant autonomy to SE tasks. It is an intuitive approach to divide high-level requirements into sub-tasks and detailed implementation, which mirrors agile and iterative methodologies (Larman, 2004) where tasks are broken down and assigned to specialized teams or individuals. By automating this process, developers are freed to focus on strategic planning, design thinking, and innovation.
  3. (3)
    Scalability to Complex Systems: The growth in complexity of software systems, with increasing lines of code, frameworks, and interdependencies, demands scalable solutions in project management and development practices. LMA systems offer an effective scaling solution by incorporating additional agents for new technologies and reallocating tasks among agents based on evolving project needs. LMA systems ensure that complex projects, which may be overwhelming for individual developers or traditional teams, can be managed effectively through distributed intelligence and collaborative agent frameworks.
Existing research has illuminated the critical roles of these collaborative agents in advancing toward the era of Software Engineering 2.0 (Lo, 2023). LMA systems are expected to significantly speed up software development and drive innovation, marking a major shift in software engineering practices. This article aims to delve deeper into the roles of LMA systems in shaping the future of software engineering, spotlighting potential applications, emerging challenges, and the vital collaboration between academia and industry to harness these advancements fully. Furthermore, we present a comprehensive research agenda, which enumerates a series of research questions to direct subsequent research efforts.
2. Preliminary
2.1. Autonomous Agent
An autonomous agent is a computational entity designed for independent and effective operation in dynamic environments (Mele, 2001). Its essential attributes are:
  • Autonomy: Independently manages its actions and internal state without external controls.
  • Perception: Detects the changes in the surrounding environment through sensory mechanisms.
  • Intelligence and Goal-Driven: Aims for specific goals using domain-specific knowledge and problem-solving abilities.
  • Social Ability: Can interact with humans or other agents, manages social relationships to achieve goals.
  • Learning Capabilities: Continuously adapts, learns, and integrates new knowledge and experiences.
2.2. LLM-based Autonomous Agent
Formally speaking, an LLM-based agent can be described by the tuple $\langle L,O,M,P,A,R\rangle$ (Cheng et al., 2024), where:
  • $L$ symbolizes the Large Language Model, serving as the agent’s cognitive core. It is equipped with extensive knowledge, potentially fine-tuned for specific domains, allowing it to make informed decisions based on observations, feedback, and rewards. Typically, an LLM suited for this role is trained on vast corpora of diverse textual data and comprises billions of parameters, such as models like ChatGPT [^1] https://openai.com/chatgpt/, Claude [^2] https://claude.ai/, and Gemini [^3] https://gemini.google.com/app. These models exhibit strong zero-shot and few-shot learning capabilities, meaning they can generalize well to new tasks with little to no additional training. Interaction with the LLM typically occurs through prompts, which guide its reasoning and responses.
  • $O$ stands for the Objective, the desired outcome or goal the agent aims to achieve. This defines the agent’s focus, driving its strategic planning and task breakdown.
  • $M$ represents Memory, which holds information on both historical and current states, as well as feedback from external interactions.
  • $P$ represents Perception, which represents the agent’s ability to sense, interpret, and understand its surroundings and inputs. This perception can involve processing structured and unstructured data from various sources such as text, visual inputs, or sensor data. Perception allows the agent to interpret the environment, transforming raw information into meaningful insights that guide decision-making and actions.
  • $A$ signifies Action, encompassing the range of executions of the agent, from utilizing tools to communicating with other agents.
  • $R$ refers to Rethink, a post-action reflective thinking process that evaluates the results and feedback, along with stored memories. Guided by this insight, the LLM-based agent then takes subsequent actions.
2.3. LLM-Based Multi-Agent Systems
A multi-agent system is a computational framework composed of multiple interacting intelligent agents that interact and collaborate to solve complex problems or achieve goals beyond the capability of any single agent (Wooldridge, 2009). These agents communicate, coordinate, and share knowledge, often bringing specialized expertise to address tasks across diverse domains.
With the integration of LLMs, LLM-Based Multi-Agent Systems have emerged. In this paper, we define that an LMA system comprises two primary components: an orchestration platform and LLM-based agents.
2.3.1. Orchestration Platform
The orchestration platform serves as the core infrastructure that manages interactions and information flow among agents. It facilitates coordination, communication, planning, and learning, ensuring efficient and coherent operation. The orchestration platform defines various key characteristics:
  1. (1)
    Coordination Models: Defines how agents interact, such as cooperative (collaborating towards shared goals) (Abdelnabi et al., 2023), competitive (pursuing individual goals that may conflict) (Wu et al., 2024), hierarchical (organized with leader-follower relationships) (Zhao et al., 2024), or mixed models.
  2. (2)
    Communication Mechanisms: Determines how the information flows between the agents: It defines the organization of communication channels, including centralized (a central agent facilitates communication (Agashe, 2023)), decentralized (agents communicate directly (Chen et al., 2023a)), or hierarchical (information flows through layers of authority (Zhao et al., 2024)). Moreover, it specifies the data exchanged among agents, often in text form. In software engineering contexts, this may include code snippets, commit messages (Zhou et al., 2023), forum posts (He et al., 2022, 2024a, 2024b), bug reports (Bettenburg et al., 2008), or vulnerability reports (Imtiaz et al., 2021).
  3. (3)
    Planning and Learning Styles: The orchestration platform specifies how planning and learning are conducted within the multi-agent system. It determines how tasks are allocated and coordinated among agents. It includes strategies like Centralized Planning, Decentralized Execution (CPDE) – planning is conducted centrally, but agents execute tasks independently, or Decentralized Planning, Decentralized Execution (DPDE) – both planning and execution are distributed among agents.
2.3.2. LLM-Based Agents
Each agent may have unique abilities and specialized roles, enhancing the system’s ability to handle diverse tasks effectively. Agents can be:
  1. (1)
    Predefined or Dynamically Generated: Agent profiles can be explicitly predefined (Hong et al., 2023) or dynamically generated by LLMs (Wang et al., 2023c), allowing for flexibility and adaptability.
  2. (2)
    Homogeneous or Heterogeneous: Agents may have identical functions (homogeneous) or diverse functions and expertise (heterogeneous).
Each LLM-based agent can be represented as a node $v_{i}$ in a graph $G(V,E)$, where edges $e_{i,j}\in E$ represent interactions between agents $v_{i}$ and $v_{j}$.
3. Literature Review
In this section, we provide a review of recent studies on LMA systems in software engineering. We categorize these applications according to the various stages of the software development lifecycle (SDLC), focusing on key areas such as requirements engineering, code generation, quality assurance, and software maintenance. Additionally, we highlight papers that explore the use of LMA systems for end-to-end software development, which encompasses multiple phases of the SDLC, rather than focusing on isolated stages.
3.1. Requirements Engineering
Requirements Engineering (Van Lamsweerde, 2000; Lo, 2024) focuses on defining and managing software system requirements. This discipline is divided into several key stages to ensure requirements meet quality standards and align with stakeholder needs. These stages include elicitation, modeling, specification, analysis, and validation (Hickey and Davis, 2004; Christel and Kang, 1992).
Elicitron (Ataei et al., 2024) is an LMA framework that focuses specifically on the elicitation stage. It utilizes LLM-based agents to represent a diverse array of simulated users. These agents engage in simulated product interactions, providing insights into user needs by articulating their actions, observations, and challenges. MARE (Jin et al., 2024) is another LMA framework that covers multiple phases of requirements engineering, including elicitation, modeling, verification, and specification. It employs five distinct agents performing nine different tasks to help generate high-quality requirements models and specifications.
3.2. Code Generation
Code generation is an essential task in software development, where source code is produced from high-level specifications, often written in natural language. In this subsection, we review the recent innovative LMA frameworks that enhance the process of code generation.
A prominent design approach in LMA systems involves role specialization and cross-validation (e.g., peer review) among agents (Li et al., 2023; Liu et al., 2023; Sun et al., 2024b; Olausson et al., 2023; Dong et al., [n. d.]). In this method, agents are assigned distinct roles, such as programmers or testers, and collaborate iteratively to enhance code quality. INTERVENOR (Wang et al., 2024b) pairs a Code Learner with a Code Teacher to iteratively improve code and fix bugs using compiler error feedback. Similarly, Self-repair (Olausson et al., 2023) involves a code agent generating programs and a feedback agent providing execution error explanations to iteratively improve code. AutoCoder (Lei et al., 2024) simulates coding tasks with a questioner and a programmer. Both agents cooperate and verify the execution through unit tests. SoA (Ishibashi and Nishimura, 2024) proposes to organize agents based on problem complexity. PairCoder (Zhang et al., 2024a) incorporates a Navigator agent for high-level planning and a Driver agent for specific implementation. The Navigator is responsible for proposing promising solution plans, selecting the current optimal plan, and directing the next iteration round based on execution feedback. The Driver follows the guidance of Navigator to undertake initial code generation, code testing, and refinement. MapCoder (Islam et al., 2024) replicates the problem-solving process in competitive programming through four specialized LLM agents. The Retrieval Agent retrieves examples of similar problems, guiding the solution process. The Planning Agent then creates a structured solution plan, assigning confidence scores to each approach. The Coding Agent follows the plan to generate and test code, while the Debugging Agent continues to correct errors by comparing the code with the initial plan.
CodeS (Zan et al., 2024) converts natural language requirements into fully functional code repositories using three specialized agents. The RepoSketcher defines the repository structure, the FileSketcher generates blueprints for each file, and the SketchFiller completes the code for each function. Cai et al. (Cai et al., 2023) present a framework to enhance human-LLM interaction through visual programming techniques. It allows users to design workflows for complex tasks using low-code interfaces, giving them more control over LLM outputs. The framework includes two agents: a Planning agent, which creates workflows, and an Executing agent, which follows the user-edited workflows. CodexGraph (Liu et al., 2024b) improves code generation, code understanding, and other software engineering tasks by navigating large code repositories. It builds graph databases using static analysis, extracting code symbols and their relationships. These symbols, represented as nodes, and their interactions, like inheritance and containment, form the edges of the graph. CodexGraph enables LLM-based agents to interact with these graph databases by generating and executing graph queries.
3.3. Software Quality Assurance
In this subsection, we review related work on testing, vulnerability detection, bug detection, and fault localization, with a focus on how LMA systems are being employed to enhance software quality assurance processes.
Testing. Fuzz4All (Xia et al., 2024) generates testing input for software systems across multiple programming languages. In this framework, a distillation agent reduces user input, while a generation agent creates and mutates inputs. AXNav (Taeb et al., 2024) is designed to automate accessibility testing. It interprets natural language test instructions and executes accessibility tests, such as VoiceOver, on iOS devices. AXNav includes a planner agent, an action agent, and an evaluation agent. WhiteFox (Yang et al., 2023a) is a fuzzing framework that tests compiler optimizations. It uses two LLM-based agents: one extracts requirements from source code, and the other generates test programs. Additionally, LMA systems are employed for tasks such as penetration testing (Deng et al., 2023), user acceptance testing (Wang et al., 2024d), and GUI testing (Yoon et al., 2024).
Vulnerability Detection. GPTLens (Hu et al., 2023) is an LMA framework for detecting vulnerabilities in smart contracts. The system includes LLM-based agents acting as auditors, each independently identifying vulnerabilities. A critic agent then reviews and ranks these vulnerabilities, filtering out false positives and prioritizing the most critical ones. MuCoLD (Mao et al., 2024) assigns roles like tester and developer to evaluate code. Through discussions and iterative assessments, the agents reach a consensus on vulnerability classification.
Bug Detection. Intelligent Code Analysis Agent (ICAA) (Fan et al., 2023) is used for bug detection in static code analysis. The agents have access to tools like web search, static analysis, and code retrieval tools. A Report Agent generates bug reports, while a False Positive Pruner Agent refines these reports to reduce false positives. Additionally, ICAA includes Code-Intention Consistency Checking, which ensures the code aligns with the developer’s intended functionality by analyzing code comments, documentation, and variable names.
Fault Localiztion. RCAgent (Wang et al., 2023b) performs root cause analysis in cloud environments by using LLM-based agents to collect system data, analyze logs, and diagnose issues. AgentFL (Qin et al., 2024) breaks down fault localization into three phases. The Comprehension Agent identifies potential fault areas, the Navigation Agent narrows down the codebase search, and the Confirmation Agent uses debugging tools to validate the faults.
3.4. Software Maintenance
In this subsection, we explore related work on debugging and code review, highlighting how LMA systems contribute to automating and improving software maintenance processes.
Debugging. Debugging involves identifying, locating, and resolving software bugs. Several frameworks, including MASAI (Arora et al., 2024), MarsCode (Liu et al., 2024a), AutoSD (Kang et al., 2023), and others (Tao et al., 2024; Ma et al., 2024; Chen et al., 2024), follow a structured process consisting of stages like bug reproduction, fault localization, patch generation, and validation. Specialized agents are typically responsible for each stage. FixAgent (Lee et al., 2024) includes a debugging agent and a program repair agent that work together to iteratively fix code by analyzing both errors and repairs. The system refines fault localization by incorporating repair feedback. The agents also articulate their thought process, improving context-aware debugging. The MASTER framework (Yang et al., 2024) employs three specialized agents. The Code Quizzer generates quiz-like questions from buggy code, the Learner proposes solutions, and the Teacher reviews and refines the Learner’s responses. AutoCodeOver (Zhang et al., 2024c) uses an agent for fault localization via spectrum-based methods, collaborating with others to refine patches using program representations like abstract syntax trees. SpecRover (Ruan et al., 2024) extends AutoCodeOver by improving program fixes through iterative searches and specification analysis, based on inferred code intent. ACFIX (Zhang et al., 2024b) targets access control vulnerabilities in smart contracts, focusing on Role-Based Access Control. It mines common RBAC patterns from over 344,000 contracts to guide agents in generating patches. DEI (Zhang et al., 2024f) resolves GitHub issues by using a meta-policy to select the best solution, integrating and re-ranking patches generated by different agents for improved issue resolution. RepoUnderstander (Ma et al., 2024) constructs a knowledge graph for a full software repository and uses Monte Carlo tree search to assist in understanding complex dependencies.
Code Review. Rasheed et al. (Rasheed et al., 2024a) developed an automated code review system that identifies bugs, detects code smells, and provides optimization suggestions to improve code quality and support developer education. This system uses four specialized agents focused on code review, bug detection, code smells, and optimization. Similarly, CodeAgent (Tang et al., 2024) performs code reviews with sub-tasks such as vulnerability detection, consistency checking, and format verification. A supervisory agent, QA-Checker, ensures the relevance and coherence of interactions between agents during the review process.
3.5. End-to-end Software Development
End-to-end software development involves managing the entire software development lifecycle, from gathering initial requirements to delivering the final product. In the context of LMA systems, each phase typically is handled by specialized agents which perform tasks suited to their expertise. The ultimate goal is to create fully autonomous systems capable of managing the entire development cycle. These systems often draw inspiration from established software development practices, such as Agile (Cohen et al., 2004), Scrum (Schwaber, 1997), and Waterfall (Petersen et al., 2009).
Pioneering works, such as ChatDev (Qian et al., 2024c) and MetaGPT (Hong et al., 2023), adopt a Waterfall model to automate the software development. ChatDev assigns agents as programmers, reviewers, and testers to handle design, coding, testing, and documentation. MetaGPT enhances the workflow by incorporating Standardized Operating Procedures (SOPs) (Belbin and Brown, 2022), with agents taking on roles like product managers and QA engineers. Recent advancements (Zhang et al., 2024d; Dong et al., 2023; Du et al., 2024) have further built on these designs, pushing the boundaries of fully autonomous development. Self-collaboration (Dong et al., 2023) presents a framework for improving code generation using multiple instances of ChatGPT. The approach simulates teamwork by assigning specific roles (analyst, coder, and tester) to distinct instances of ChatGPT. CTC (Du et al., 2024) involves multiple teams of LLM-based agents working together to generate high-quality software. Each team independently proposes solutions, and cross-communication allows them to refine their outputs by eliminating suboptimal proposals.
AgileCoder (Nguyen et al., 2024) and AgileGen (Zhang et al., 2024e) apply Agile methodologies in software development. AgileCoder (Nguyen et al., 2024) mirrors Agile’s role-based collaboration, assigning tasks like Product Manager and Developer to work in sprints. It also features a Dynamic Code Graph Generator, which updates code dependencies in real time, supporting context-aware generation and refinement. AgileGen (Zhang et al., 2024e) combines human-AI collaboration with Agile principles, using Gherkin for user acceptance criteria, ensuring alignment between requirements and code, and reducing user effort.
TGen (Mathews and Nagappan, 2024) explores Test-Driven Development (TDD), where test cases guide GPT-4 to generate correct code. LCG (Lin et al., 2024a) emulates various software process models, including Waterfall, TDD, and Scrum.
Additionally, several works (Rasheed et al., 2024b; Lin et al., 2024b; Qian et al., 2024b; Sami et al., 2024) adopt iterative and adaptive methodologies without explicitly detailing their models. For instance, ToP (Lin et al., 2024b) utilizes process models to guide agents in software development, which are refined through heuristic algorithms and process mining (Van Der Aalst, 2012). Qian et al. (Qian et al., 2024b) propose an iterative experience refinement framework that enables agents to continuously adapt to software development by leveraging prior experiences. This framework supports ongoing learning through the acquisition, utilization, and selective elimination of experiences. Co-Learning (Qian et al., 2024a) fosters collaboration through co-tracking, co-memorizing, and co-reasoning modules among agents. GameGPT (Chen et al., 2023b) automates and streamlines game development, employing various roles like game designers and code reviewers to enhance decision-making in game creation.
4. Research Agenda
Previous research has laid the groundwork for the exploration of LMA systems in software engineering, yet this domain remains in its nascent stages, with many critical challenges awaiting resolution. In this section, we outline our perspective on these challenges and suggest research questions that could advance this burgeoning field. We envision two phases for the development of LMA systems in software engineering. We discuss each of these phases below and suggest a series of research questions that could form the basis of future research projects.
4.1. Phase 1: Enhancing Individual Agent Capabilities
Indeed, the effectiveness of an LMA system is closely linked to the capabilities of its individual agents. This first phase is dedicated to improving these agents’ skills, with a particular focus on adaptability and the acquisition of specialized skills in SE. The potential of individual LLM-based agents in SE is further explored through our initial research questions:
  1. (1)
    What SE roles are suitable for LLM-based agents to play and how can their abilities be enhanced to represent these roles?
  2. (2)
    How to design an effective, flexible, and robust prompting language that enhances LLM-based agents’ capabilities?
4.1.1. Refining Role-Playing Capabilities in Software Engineering
The role-playing capabilities of LLM-based agents are pivotal within an LMA system (Wang et al., 2023d). Given the complex nature of SE tasks, we require specialized agents to focus on specific information, apply specialized knowledge, and think from diverse perspectives to generate meaningful and effective outputs. While LLM-based agents have shown proficiency in simulating well-recognized roles like generic software developers or project managers (Hong et al., 2023; Qian et al., 2024c), they often struggle to accurately represent specialized expert roles that demand extensive domain expertise. For example, roles involving vulnerability detection or security auditing require a nuanced understanding of security protocols, threat modeling, and the latest vulnerabilities, which goes beyond general programming knowledge. Similarly, roles like database optimization specialists or embedded systems engineers demand specialized skills and insights that are challenging for LLMs to emulate accurately. This shortfall arises primarily because they are trained on general text corpora that lack the specialized content necessary to grasp the nuanced expertise required in SE (Hou et al., 2023). This limitation highlights the need for integrating domain-specific expertise into LLMs to better support specialized roles within software engineering. To address this limitation, we propose a structured and actionable three-step approach encompassing the identification, assessment, and enhancement of role-playing abilities, which are:
Step 1: Identifying and Prioritizing Key SE Roles.
Step 2: Assessing LLM-Based Agents’ Competencies Against Role Requirements.
Step 3: Enhancing Role-Playing Abilities Through Targeted Training.
The first step focuses on identifying key SE roles, prioritizing those with high industry demand and the potential to substantially boost productivity. This involves:
  1. (1)
    Market Analysis: To begin, we embark on a comprehensive market analysis. It is crucial to assess not only the current trends and needs within the SE sector but also to anticipate future shifts influenced by the integration of LLM-based agents. This analysis involves leveraging various resources such as market reports, job postings, industry forecasts, and technology trend analyses. Platforms like LinkedIn Talent Insights [^4] https://www.linkedin.com/products/linkedin-talent-insights/, Gartner reports [^5] https://www.gartner.com/en/products/special-reports, and Stack Overflow Developer Surveys [^6] https://survey.stackoverflow.co/2024/ may also offer valuable data to inform this assessment. The focus should be on identifying roles that are in high demand and demonstrate rapid growth, especially those requiring specialized skills not typically found among generalist developers. For example, machine learning engineers or cloud architects. Additionally, positions where LLM-based agents could significantly enhance productivity, reduce costs, or accelerate innovation should be evaluated. A key component of this analysis should be determining whether the market has already begun shifting away from recruiting humans for tasks that LLM-based agents can perform, such as routine coding or simple bug fixing. Identifying these trends will help distinguish between roles that are still in demand and those where LLMs have reduced the need for human expertise.
  2. (2)
    Stakeholder Engagement: Engaging comprehensively with a diverse group of stakeholders is essential. This process validates the findings from the market analysis and ensures that the selected roles align with real-world needs. It involves consulting industry professionals who have hands-on experience in the identified roles. This engagement can provide practical insights and challenges associated with these positions. Collaboration with HR departments from leading technology companies is also important. It helps gather perspectives on current hiring trends, skill shortages, and the most sought-after competencies. Additionally, academic experts and researchers can offer forward-thinking views on emerging technologies and methodologies. By incorporating feedback from these various sources, the selection of key roles becomes more robust to reflect both current industry demands and future directions.
  3. (3)
    Value Addition Modeling: The next crucial step is value addition modeling (Mendes et al., 2018), which evaluates the potential advantages that LLM-based agents could bring to each prioritized role. This process involves constructing detailed, data-driven models to analyze key performance indicators such as efficiency improvements, cost reductions, quality enhancements, and the acceleration of innovation resulting from the integration of agents. Pilot projects can be deployed to gather empirical data on these metrics when LLM-based agents are applied to specific tasks. Important factors to consider include the automation of repetitive tasks, the augmentation of human capabilities, and the enablement of new functionalities that were previously unattainable. It is important to note that the value added by LLM-based agents can differ significantly across different domains; for example, roles in software development may prioritize automation, whereas domains like systems architecture might see more value in LLMs augmenting complex decision-making around resource allocation or performance optimization, where human expertise and contextual understanding remain essential. By quantifying these value propositions, organizations can allocate resources more strategically to roles where LLM-based agents are likely to yield the highest return on investment.
The second step involves understanding the limitations of LLM-based agents relative to the demands of the identified SE roles:
  1. (1)
    Competency Mapping: Competency mapping (Kaur and Kumar, 2013) entails developing comprehensive competency frameworks for each specialized role. These frameworks define the essential skills, knowledge areas, and competencies required, encompassing both technical and soft skills. For instance, technical skills might encompass proficiency in specific programming languages, tools, methodologies, and domain-specific knowledge. For a machine learning engineer, this would include expertise in algorithms, data preprocessing, model training, and tools such as TensorFlow [^7] https://www.tensorflow.org/ or PyTorch [^8] https://pytorch.org/. Soft skills include skills like problem-solving, critical thinking, and collaboration. Clearly outlining these competencies creates a benchmark against which the agents’ abilities can be measured.
  2. (2)
    Performance Evaluation: The next phase is performance evaluation, which involves designing or selecting tasks that closely replicate the real-world challenges associated with each role. These tasks should be practical and scenario-based to accurately gauge the agents’ capabilities. They should assess a wide range of competencies, from technical execution to critical thinking. For example, in evaluating a DevOps engineer, the agent might be tasked with automating a deployment pipeline using tools like Jenkins [^9] https://www.jenkins.io/ or Docker [^10] https://www.docker.com/, or troubleshooting a continuous integration failure. Such tasks allow for a thorough assessment of both technical and soft skills.
  3. (3)
    Gap Analysis: This step compares the agents’ outputs with the expected outcomes for each task. Key areas where the agents underperform–such as misunderstanding domain-specific terminology, neglecting security best practices, or failing to optimize code–are identified and documented. This analysis emphasizes both the agents’ strengths and weaknesses, offering valuable insights into recurring patterns of errors or misconceptions.
  4. (4)
    Expert Consultation and Iterative Refinement: To further refine the evaluation process, expert consultation and iterative refinement are essential. By engaging with SE professionals who specialize in the assessed roles, qualitative feedback on the agent’s performance can be obtained. These experts provide insights into subtle nuances that may not be captured through quantitative metrics. For instance, while the agent’s code may work, it might not follow best practices or address scalability. This feedback helps refine evaluation methods, update competency frameworks, and uncover deeper issues in the agent’s understanding.
The final step involves tailoring the LLM-based agents to effectively represent the identified SE roles through specialized training and prompt engineering. :
  1. (1)
    Curating Specialized Training Data: At first, this involves creating training datasets that reflect the unique requirements of each specific role. A comprehensive corpus should be built from a variety of sources, including technical documentation such as API guides, technical manuals, and user guides to provide in-depth knowledge of specific technologies. It is also important to incorporate academic and industry research papers, case studies, and whitepapers to capture the latest developments, best practices, and theoretical foundations. Additionally, discussions from forums and software Q&A sites like Stack Overflow [^11] https://stackoverflow.com/, Reddit [^12] https://www.reddit.com/, and specialized industry forums can provide practical problem-solving approaches and real-world challenges faced by professionals.
  2. (2)
    Fine-tuning the LLM: After preparing the data, the curated datasets are used to fine-tune the LLM-based agents. Advanced techniques like parameter-efficient fine-tuning (PEFT) (Liu et al., 2022) are often employed to optimize both efficiency and accuracy.
  3. (3)
    Designing Customized Prompts: A key step is designing prompts tailored to improve the agents’ role adaptability. These prompts should clearly define the role, tasks, and goals to ensure the agent understands the requirements. For instance, in a cybersecurity analyst role, the prompt should outline specific security protocols, potential vulnerabilities, and compliance standards. Contextual instructions, including relevant background, constraints, and examples, help the agent grasp task nuances. Creating a library of effective prompts for various scenarios can also serve as reusable templates for future tasks.
  4. (4)
    Continuous Learning and Adaptation: To keep agents aligned with industry developments, continuous adaptation mechanisms are essential. Training data should be regularly updated, and models may be retrained to incorporate new technologies, best practices, and trends in software engineering. Monitoring systems can track agent performance over time, enabling proactive adjustments and continuous improvement. Additionally, agents should be guided to consistently reference the latest documentation and standards to ensure their outputs remain relevant and accurate.
While LMA roles may overlap with traditional software engineering roles, it is important to recognize that they are not necessarily the same, as LMA roles often involve specialized, collaborative tasks suited for agent-based systems. By systematically identifying key roles, assessing agent competencies, and enhancing their capabilities through targeted fine-tuning, we aim to significantly improve the effectiveness of LLM-based agents in specialized SE roles.
4.1.2. Advancing Prompts through Agent-oriented Programming Paradigms
In previous sections, we discussed the benefits of well-crafted prompts. However, crafting such prompts is indeed challenging due to the need for a versatile, effective, and robust framework applicable across various scenarios. Natural language, while flexible, often contains ambiguities and inconsistencies that LLMs may misinterpret. This highlights the need for a specialized prompting language designed to augment the cognitive functions of LLM-based agents. Natural language is inherently designed for human communication, relying on shared context and intuition that LLMs lack. LLMs interpret text based on statistical patterns from large datasets, which can lead to different interpretations than those intended for humans (Sun et al., 2024a; Zeng et al., 2022). To bridge this gap, we propose the development of a formalized prompting language that treats LLMs as the primary audience. Such a language aims to minimize ambiguities and ensure clear, deterministic instructions, resulting in more reliable and accurate outputs.
Agent-oriented programming (AOP) (Shoham, 1993) offers a promising foundation for this approach. Similar to how Object-Oriented Programming (OOP) (Wegner, 1990) organize objects, AOP treats agents as fundamental units, focusing on their reasoning, objectives, and interactions. An AOP-based prompting language could enable the precise expression of complex tasks and constraints, allowing LLM-based agents to perform their roles with greater efficiency and accuracy. By clearly defining instructions within the AOP framework, we can reduce ambiguity, mitigate hallucinations, and enhance task execution.
However, creating a new language for prompts presents its own challenges. Such a language must be expressive enough to handle diverse and complex tasks, yet simple enough for users to easily adopt. Balancing functionality and usability will be critical to its success. Additionally, this process may involve tailoring prompts specifically for different LLM models and their versions, as variations in model architectures, training data, and capabilities can affect how they interpret and respond to prompts. What works effectively for one model may not perform as well for another, necessitating careful adjustments. While AOP-based prompting may not be the final solution, it represents an important step toward developing an AI-oriented language with grammar tailored specifically for LLMs. This new approach could further refine communication with LLM-based agents, reducing misinterpretation and significantly enhancing overall performance.
4.2. Phase Two: Optimizing Agent Synergy
In Phase Two, the spotlight turns towards optimizing agent synergy, underscoring the importance of collaboration and how to leverage the diverse strengths of individual agents. This phase delves into both the internal dynamics among agents and the role of external human intervention in enhancing the efficacy of the LMA system. Key research questions guiding this phase include:
  1. (3)
    How to best allocate tasks between humans and LLM-based agents?
  2. (4)
    How can we quantify the impact of agent collaboration on overall task performance and outcome quality?
  3. (5)
    How to scale LMA systems for large-scale projects?
  4. (6)
    What industrial organization mechanisms can be applied to LMA systems?
  5. (7)
    What strategies allow LMA systems to dynamically adjust their approach?
  6. (8)
    How to ensure security among private data sharing within LMA systems?
4.2.1. Human-Agent Collaboration
Optimally distributing tasks between humans and LMA agents to leverage their respective strengths is essential. Humans bring unparalleled creativity, critical thinking, ethical judgment, and domain-specific knowledge (Markauskaite et al., 2022). In contrast, LMA agents excel at rapidly processing large datasets, performing repetitive tasks with high accuracy, and detecting patterns that might elude human observers. Aligning LMA systems with human intentions and expertise is critical, but integrating human involvement throughout the software development lifecycle remains insufficiently explored. Identifying the most effective roles for humans, optimizing feedback mechanisms, and determining the appropriate moments for human intervention present significant challenges.
To facilitate seamless collaboration, designing intuitive user-friendly interfaces and interaction protocols is essential (Wang et al., 2024a). Natural language interfaces and adaptive visualization techniques can make interactions more accessible. These interfaces should efficiently present agent outputs in a digestible format and collect user feedback, while also managing the cognitive load on human collaborators. It is important to note that these interfaces may need to be tailored differently for each human role, as the needs of a project manager, a software developer, and a quality assurance engineer will vary significantly. Given the complexity of information generated during the agents’ workflows (Josifoski et al., 2023), designing such interfaces poses challenges. For instance, presenting modifications suggested by an agent at varying levels of abstraction ensures that each stakeholder can engage with the information at the right depth. A project manager might focus on the broader implications, such as the high-level impact on project timelines or deliverables, whereas a developer or architect might drill down into specific implementation details. Role-specific interfaces will be key to ensuring each stakeholder can effectively collaborate with the agents and extract the necessary information in a manner suited to their specific responsibilities.
Additionally, developing predictive models to determine the optimal human-to-agent ratio across different project types and stages is a fundamental concern. These models must assess factors such as project complexity, time constraints, project priorities, and the specific capabilities and limitations of both human participants and LMA agents. By doing so, tasks can be allocated in a manner that fully harnesses both human ingenuity and agent efficiency throughout the project. Machine learning techniques could also be leveraged to analyze historical project data to predict effective collaboration strategies.
4.2.2. Evaluating the LMA systems
Recently, numerous complex benchmarks have been proposed to challenge the capabilities of LLMs in critical aspects of software development, such as code generation (Zhuo et al., 2024; Jain et al., 2024; Liu et al., 2024c). While these benchmarks have advanced the field by providing measurable metrics for individual tasks, their focus on isolated problem-solving reveals limitations as software engineering projects become more complex.
Software engineering is inherently collaborative, with key activities like joint requirements gathering, code integration, and peer reviews playing essential roles in the process. Current benchmarks often overlook these aspects, failing to assess how well LLMs perform in tasks that require cooperation and collective decision-making. There is a growing need for benchmarks that evaluate the cooperative abilities of LLMs in multi-agent settings, particularly for software engineering tasks. These benchmarks should simulate real-world collaborative scenarios where LLM agents work together to achieve common development goals.
Such benchmarks should include tasks where agents must:
  1. (1)
    Participate in collaborative design: Agents should contribute ideas, propose design solutions, and converge on a unified architecture that balances trade-offs.
  2. (2)
    Delegate and coordinate tasks: Effective task division is crucial. Agents should assign responsibilities based on expertise, manage dependencies, and adjust as the project evolves.
  3. (3)
    Identify conflicts and negotiate: In collaborative settings, disagreements are inevitable. First, LLMs often struggle to identify conflicts in real time unless explicitly guided to do so (Abdelnabi et al., 2023). Therefore, agents should be evaluated on their ability to recognize these conflicts—whether in logic, goals, or execution. Moreover, agents should be tested on their ability to handle conflicts constructively. This includes proposing compromises, engaging in constructive negotiation, and ensuring that the team remains aligned with the overarching objectives. Evaluations should focus on the agents’ capacity to balance competing priorities, mitigate misunderstandings, and foster consensus, all while maintaining progress toward shared goals.
  4. (4)
    Integrate components and perform peer reviews: Agents should seamlessly integrate their work, review each other’s code for quality assurance, and provide constructive feedback.
  5. (5)
    Proactive Clarification Request : Agents should not assume complete understanding when uncertainty arises. Instead, they should preemptively ask for additional information or clarification to avoid potential errors or misunderstandings. Evaluating agents on this ability ensures they are capable of identifying gaps in their knowledge or instructions and can actively seek out the necessary context or data to complete tasks effectively.
To develop such benchmarks, we need to create realistic project scenarios that require multi-agent collaboration over extended periods. These scenarios should reflect common software development challenges, such as evolving requirements and tight deadlines. Additionally, platforms or sandboxes must be built to provide controlled environments where collaborative interactions between agents can be observed and measured. These platforms should establish clear interaction rules, including languages, formats, and communication channels, to facilitate effective information exchange.
Most importantly, comprehensive metrics must be developed to assess not just the final output, but also the collaboration process itself. These metrics could measure communication efficiency, ambiguity resolution, conflict management, adherence to best practices, and overall project success.
4.2.3. Scaling Up for Complex Projects.
As software projects become more complex, a single LLM-based agent may hit its performance limits. Inspired by the scaling properties of neural models (Bahri et al., 2024), LMA systems can potentially enhance performance by increasing the number of agents within the system. While adding more agents can provide some benefits, handling more complex projects introduces challenges that require more refined solutions.
First, as software projects grow in size and complexity, breaking down high-level requirements into manageable sub-tasks becomes more difficult. It is not just about handling more tasks but also managing the intricate interdependencies between them. A hierarchical task decomposition approach can help, where higher-level agents oversee broader objectives and delegate specific tasks to lower-level agents. This structure streamlines planning and makes global task allocation more efficient.
Second, as the number of agents increases, so does the complexity of communication. Coordinating multiple agents can lead to communication bottlenecks and information overload. Additionally, large-scale software projects challenge the memory capacity of individual agents, making it harder to store and process the extensive information required. Efficient communication protocols and message prioritization are crucial to mitigating these issues. For instance, agents can use summarized updates instead of detailed reports, reducing communication overhead and memory usage. From the outset, the system should be designed with scalability in mind, ensuring that both software and hardware resources can expand efficiently as the number of agents increases.
Moreover, with more agents comes the risk of inconsistencies and conflicts in the shared information. A centralized knowledge repository or shared blackboard system can ensure that all agents have access to consistent, up-to-date information, acting as a single source of truth and minimizing the spread of misinformation. Robust error handling mechanisms should also be implemented to detect and correct issues autonomously before they escalate into significant failures.
Finally, as the number of agents grows, so do the rounds of discussion and decision-making, which can slow down progress. To avoid this, decision-making hierarchies or consensus algorithms can streamline the process. For example, only a subset of agents responsible for a specific module may need to reach a consensus, rather than involving the entire agent network.
4.2.4. Leveraging Industry Principles
As LLM-based agents can closely mimic human developers in SE tasks, they can greatly benefit from adopting established industry principles and management strategies. By emulating organizational frameworks used by successful companies, LMA systems can improve their design and optimization processes. These industrial mechanisms enable LMA systems to remain agile, efficient, and effective, even as project complexities grow.
Leadership and governance structures from industrial organizations provide valuable insights for designing LMA systems. Project management tools and practices, essential for coordinating large development teams, can be applied to LMA systems to enhance their operational efficiency. Using established project management frameworks, LMA systems can monitor progress, allocate resources, and manage timelines effectively. Agents can dynamically update task boards, report milestones, and adjust workloads in real-time based on project data. This not only improves transparency but also allows for early detection of bottlenecks or delays, ensuring projects stay on track.
Incorporating design patterns and software architecture best practices further strengthens LMA systems (Lee et al., 2024). By adhering to these principles, agents can produce well-structured, maintainable code that is scalable and reusable. This reduces technical debt and ensures that the solutions developed by LMA systems are easier to integrate, maintain, and expand in the future.
4.2.5. Dynamic Adaptation.
In the context of software development, predicting the optimal configuration for LMA systems at the outset is unrealistic due to the inherent complexity and variability of tasks (Leffingwell and Widrig, 2000). The dynamic nature of software requirements and the unpredictable challenges that arise during development necessitate systems that can adapt on the fly (Liu et al., 2023). Therefore, LMA systems must be capable of dynamically adjusting their scale, strategies, and structures throughout the development process.
To minimize redundant work, LMA systems should continuously evaluate existing solutions (Wang et al., 2024c), identifying reusable elements for new requirements. By learning from each development cycle, the system can recognize patterns of efficiency and inefficiency, enabling it to make informed decisions when handling similar tasks in the future or adapting existing solutions to new requirements.
A key element of dynamic adaptation is the ability to automatically adjust the number of agents involved in a project (Guo et al., 2024). This includes not only scaling the number of agents up or down as needed but also generating new agents with new specialized roles to meet emerging task requirements, ensuring both efficiency and responsiveness. Additionally, the system can replicate agents in existing roles to manage increased workloads. Furthermore, LMA systems can generate new agents that come equipped with contextual knowledge of the project—such as its history, current state, and objectives—by accessing shared knowledge bases, project documentation, and recent communications. This allows new agents to integrate smoothly and contribute effectively right from the start, reducing onboarding time and minimizing disruptions.
Another key component is the dynamic redefinition of agent roles (Jouvin and Hassas, 2002). As the project evolves, certain roles may become obsolete while new ones emerge. LMA systems should be capable of reassigning roles to agents or modifying their responsibilities to better align with current project needs. This flexibility enhances the system’s ability to adapt to changing requirements and priorities.
Dynamic adaptation also involves the reallocation of memory and computational resources. As agents are added or removed and tasks shift in complexity, the system must efficiently distribute resources to where they are most needed. This may include scaling computational power for agents handling intensive tasks or increasing memory allocation for agents processing large datasets. Effective resource management ensures that the system operates optimally without unnecessary strain on infrastructure.
Finally, the uncertainty of the software development process makes it challenging to define effective termination conditions (Smith and Merritt, 2020). Relying solely on predefined criteria may result in infinite loops or premature task completion. To address this, LMA systems must incorporate real-time monitoring and feedback loops to continuously evaluate progress. Machine learning techniques can help predict optimal stopping points by analyzing historical data and current performance metrics, allowing for informed adjustments to task completion criteria as the project evolves.
4.2.6. Privacy and Partial Information.
In multi-organizational software development projects, data often resides in silos due to privacy concerns, proprietary restrictions, and regulatory compliance requirements (Paasivaara et al., 2008). Each entity may have its own data governance policies and competitive considerations that limit data sharing. This fragmentation poses significant challenges in enabling agents to access necessary information while ensuring that privacy is maintained. Moreover, a lack of transparency in data sources and processes can exacerbate the risk of privacy violations, which may go unnoticed if data handling activities are not fully visible to all parties (Crawford and Schultz, 2014).
To address these challenges, robust and fine-grained access control mechanisms must be implemented across organizational boundaries. It is essential to prevent unauthorized access while still accommodating the varied data access needs of the system. Traditional models like Role-Based Access Control (RBAC) (Sandhu, 1998) and Attribute-Based Access Control (ABAC) (Hu et al., 2015) may need to be extended to handle the dynamic nature of multi-agent systems effectively. Establishing protocols that allow agents to share insights derived from sensitive data, without exposing the data itself, is critical. Advanced privacy-preserving techniques like Differential Privacy (Dwork, 2006), Secure Multi-Party Computation (SMPC) (Goldreich, 1998), Federated Learning (Kairouz et al., 2021), or Homomorphic Encryption (Yi et al., 2014) can be leveraged to ensure that agents collaborate without compromising data privacy.
Moreover, compliance with data protection laws such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017) in the EU and the California Consumer Privacy Act (CCPA) (Pardau, 2018) in the U.S. is crucial. LMA systems should follow privacy-by-design principles, ensuring that data subjects’ rights are upheld, and that data processing activities remain transparent and lawful. This includes implementing mechanisms for data minimization, consent management, and honoring the right to be forgotten.
For non-sensitive data, integrated data storage solutions can reduce redundancy, improve data consistency, and increase efficiency. This can be achieved through distributed databases accessible to authorized agents, along with data synchronization mechanisms to ensure agents have up-to-date information in real time. Additionally, using technologies like blockchain (Zheng et al., 2018) and distributed ledgers (Sunyaev and Sunyaev, 2020) can enhance transparency, traceability, and tamper-resistance in recording agent transactions and data access events, fostering greater trust among collaborating entities.
5. Conclusion and Future Work
In conclusion, this paper presents a roadmap for advancing LLM-based Multi-Agent systems in software engineering, outlining immediate and long-term research priorities. As part of this vision, we conducted a review of recent works in software engineering that apply LMA systems across various stages of the software development lifecycle. This review highlights key contributions in areas such as requirements engineering, code generation, software quality assurance, and maintenance. We further outline a structured research agenda to advance LMA integration in SE. Future work will concentrate on addressing the set of critical research questions identified, with a focus on enhancing LMA capabilities and optimizing their synergy within software development processes.
In the immediate term, efforts will be dedicated to enhancing the capabilities of LLM-based agents in representing specialized SE roles. This will involve creating specialized datasets and pre-training tasks that mirror the complex realities of SE tasks. Additionally, there will be a focus on formulating advanced prompting strategies, which can refine the agents’ cognitive functions and decision-making skills. Looking towards the longer-term objectives, the emphasis will transition towards optimizing the synergy between agents. The initial step involves examining optimal strategies for task allocation between humans and LLM-based agents, capitalizing on the unique strengths of both entities. Further, we need to delve into developing scalable methodologies for LMA systems, enabling LMA systems to efficiently orchestrate and complete large-scale, multifaceted SE projects. Furthermore, ensuring the privacy and confidentiality of data within LMA systems is critical. This entails exploring data management and access control mechanisms to protect sensitive information while still enabling the essential exchange of insights among project stakeholders. By systematically exploring these research questions, our goal is to drive innovation in LMA systems for software engineering, aiming to create a more cohesive, effective, and flexible LMA-driven development process.